# FCS-HGNN: Flexible Multi-type Community Search in Heterogeneous Information Networks

### v2

@@ 4. FCS-HGNN
In this section, we first provide an overview of the FCS-HGNN, followed by a detailed description of each component in FCS-HGNN.
4.1. Framework
@@
We illustrate the overall framework of FCS-HGNN in Figure 2, which includes three main steps: Heterogeneous and Query Information Preprocessing, Information Integration, and Search Algorithm. In heterogeneous and query information preprocessing, since different types of nodes in HINs usually have different semantic features, we design a type-shared feature projection to map different types of features to a unified feature space, thereby obtaining a new graph denoted as $\mathcal{H}_{h}$. In addition, we incorporate query-dependent information into HIN, denoted $\mathcal{H}_{q}$, where the graph structure maintains consistency with $\mathcal{H}$. To better utilize information from different views provided by $\mathcal{H}_{h}$ and $\mathcal{H}_{q}$, we design two encoders in information integration: the heterogeneous encoder (abbreviated as ”Hetero. Encoder” in Figure 2) and the query encoder. In each encoder, we propose an edge semantic attention mechanism to selectively aggregate neighbor information based on different types of relations, thereby capturing more fine-grained heterogeneous information. Meanwhile, FCS-HGNN integrates complementary information (query-dependent and query-independent information) after each layer using an attention mechanism. Ultimately, FCS-HGNN utilizes this integrated information to predict the probability of each node belonging to the community. In the search algorithm, we search the target community by BFS based on the probability of each node and user-specified node types $\mathcal{S}_{\mathcal{A}}$.
@@ 4.2. Heterogeneous & Query Information Preprocessing
In the community search task over HINs, we primarily have to deal with two aspects of information: the query-dependent information provided by the user and the query-independent information inherent in HINs, such as the node features and the graph structure. Differing from homogeneous graphs, HINs typically comprise multiple types of nodes and edges, presenting significant challenges to community search due to their heterogeneity. Existing studies on community search usually deal with the heterogeneity among nodes through predefined meta-paths (Fang et al., 2020; Qiao et al., 2021; Jiang et al., 2022a; Zhou et al., 2023; Liu et al., 2023) or user-specified relationship constraints (Jian et al., 2020). However, metapath-based methods are designed to identify single-type communities with nodes of the same type rather than multi-type communities. On the other hand, constraint-based methods require users to have a thorough understanding of community patterns; otherwise, they may fail to identify any community.
@@ 4.2.1. Type-shared Feature Projection
To address these issues, we propose a type-shared feature projection to elegantly deal with the heterogeneity. In HINs, nodes of different types typically reside in distinct feature spaces, which hinders the information aggregation among them (Wang et al., 2019; Yang et al., 2023). Therefore, we assume the existence of a unified feature space $\mathcal{U}\in\mathbb{R}^{d_{u}}$. For each node type $t\in\mathcal{T}$, where $\mathcal{T}=\{\psi(v_{i})|v_{i}\in V\}$, we project the original feature of each node $x[v_{i}]$ to the unified feature space $\mathcal{U}$ through a learnable type-shared linear layer with the following formula:
@@
where $\forall v_{i}\in V$, $t$ represents the type of node $v_{i}$, i.e., $\psi(v_{i})=t$, $W_{t}\in\mathbb{R}^{d_{t}\times d_{u}}$ is a learnable and type-specific linear layer, $d_{t}$ denotes the feature dimension of the original space for node type $t$, $d_{u}$ denotes the feature dimension of the unified space $\mathcal{U}$.
Equation 1 enables nodes of different types to effectively aggregate information in a unified space, alleviating the burden on systems or users as they no longer need to design predefined meta-paths or a series of relationship constraints to handle heterogeneity. It is worth noting that we still preserve the heterogeneity of edges in $\mathcal{H}_{h}$ rather than disregarding it, and propose an edge semantic attention mechanism (elaborated in Section 4.3) to capture more fine-grained heterogeneous information.
@@ 4.2.2. Query-dependent Graph
In addition to the node feature and graph structure inherent in HINs, query information is also very crucial for community search. Therefore, we encode query-dependent information $\mathcal{X}_{q}$ into the HIN $\mathcal{H}$, denoted as $\mathcal{H}_{q}$, where the graph structure maintains consistency with $\mathcal{H}$. Here, $\mathcal{X}_{q}\in\mathbb{R}^{n\times 1}$ is associated with the query node $q$ and is composed of $x_{q}[v_{i}]$, $\forall{v_{i}}\in V$.
@@
The query information $\mathcal{X}_{q}$ shares the same semantics, indicating whether they are query nodes, allowing nodes of different types to reside in the same feature space. However, the same as $\mathcal{H}_{h}$, we maintain the edge heterogeneity in $\mathcal{H}_{q}$, preserving the type of each edge. This is done to effectively preserve heterogeneous structural information around the query node during the information integration, rather than simply treating $\mathcal{H}_{q}$ as a homogeneous graph.
As shown in Figure 2, we obtain $\mathcal{H}_{h}$ and $\mathcal{H}_{q}$ from $\mathcal{H}$ by employing type-shared feature projection and introducing query-dependent information. $\mathcal{H}_{h}$ and $\mathcal{H}_{q}$ offer two distinct views containing complementary information. The former describes the features of nodes with various types in a unified space, which is beneficial to information extraction and integration in the subsequent steps. The latter delineates the local structural information around the query node, which is crucial for community search.
@@ 4.3.1. Edge Semantic Attention Mechanism
Different from homogeneous graphs, edges in HINs contain complex relationships, which are crucial for understanding the relationships between nodes and mining valuable community information. For example, in bibliography networks (Shi et al., 2012a; Sun et al., 2011), author-paper (A-P) edges represent authoring relationships, and paper-venues (P-V) edges represent publishing relationships. Therefore, the heterogeneity inherent in edges still requires careful consideration. Existing methods (Fang et al., 2020; Jian et al., 2020; Jiang et al., 2022a; Zhou et al., 2023; Liu et al., 2023) treat different types of relations equally when identifying communities. However, in multi-type communities, the contributions of different relations usually vary across different communities.
To address the above issues, we propose the edge semantic attention mechanism as follows:
@@
In real-world heterogeneous information networks, not all networks have edge features. Therefore, we consider the following two scenarios:
  1. (1)
    The heterogeneous network includes edge features;
  2. (2)
    The heterogeneous network lacks edge features.
In the first scenario, where the heterogeneous information network encapsulates edge features, we can seamlessly apply equation 3. In the second scenario, to encode the information of edge types, we introduce an embedding matrix $\mathbf{E}$, which assigns a learnable feature vector to each edge type. Given =$\mathcal{E}=\{\phi(e_{i,j})|e_{i,j}\in E\}$ is the set of edge types in the graph $\mathcal{H}$, $\mathcal{E}_{t}$ denotes the $t$-th element in $\mathcal{E}$ and $\mathcal{E}_{t}=\phi(e_{i,j})$, we formalize the embedding matrix as follows:
(4) $\vec{\mathbf{e}}_{i,j}=\mathbf{E}[t,:]$
where $t\in\{1,2,\ldots,|\mathcal{E}|\}$ is a subscript of $\mathcal{E}$ indicating the corresponding type of edge $e_{i,j}$, $\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d_{e}}$, $|\mathcal{E}|$ denotes the number of edge types, $d_{e}$ denotes the feature dimension of edges. Through this approach, even in the absence of edge features, we are able to capture and utilize the intrinsic heterogeneity of the edges. This not only ensures the richness of the aggregate information but also enhances the adaptability of our algorithm to diverse HINs, thereby improving the overall effectiveness of multi-type community search.
@@ 4.3.2. Heterogeneous Encoder
The heterogeneous encoder is dedicated to capturing query-independent information, including node features as well as structural information in $\mathcal{H}_{h}$. Compared to previous studies in HINs (Jian et al., 2020; Fang et al., 2020; Jiang et al., 2022a; Zhou et al., 2023), the heterogeneous encoder dynamically considers the contributions of different relations to communities, rather than treating them equally.
The heterogeneous encoder is based on the idea of graph neural networks (GNN) (Zhang et al., 2019; Lv et al., 2021), equipped with the edge semantic attention mechanism, and iteratively combines the neighbor information of each node as follows:
@@
Through Equation 5, we can uniformly consider structural information and node features with any dimension, avoiding ignoring node features or only considering low-dimensional features as in previous studies (Jian et al., 2020; Fang et al., 2021; Zhou et al., 2023). Compared to the vanilla GNN-based community search methods (Gao et al., 2021; Jiang et al., 2022b; Li et al., 2023) in homogeneous graphs, our proposed heterogeneous encoder integrates edge types and dynamically considers the contributions of different relations to communities, which significantly enhances the understanding of complex relationships in HINs.
4.3.3. Query Encoder
The query encoder is dedicated to obtaining query-dependent information in $\mathcal{H}_{q}$, including the local structural information around the query node, due to the query-dependent nature of community search. Compared to previous studies (Jian et al., 2020; Jiang et al., 2022b; Zhou et al., 2023), our proposed query encoder aggregates query information based on different types of edges to promote the model’s understanding of complex relationships in HINs. Specifically, the query encoder consists of multiple GNN layers, each incorporating the edge semantic attention mechanism as follows:
[…]
where $\vec{\mathbf{h}}_{qi}^{l}$ denotes the query-dependent representation of node $v_{i}$ in the $l$-th layer, and $\vec{\mathbf{h}}_{qi}^{0}=x_{q}[v_{i}]$. Note that for brevity, the learnable parameters, such as $\alpha_{ijk}^{l}$, $\mathbf{W}_{k}^{l}$ and $\mathbf{W_{r}}^{l}$, are no longer distinguished in Equations 5 and 6, but they train these parameters separately. The difference between Equation 5 and Equation 6 lies in their focus on different views, dealing with query-independent and query-dependent information, respectively. Through query encoder, we can better capture the local structural information around query nodes to facilitate community search.
@@ 4.3.4. Feature Fusion
The above heterogeneous encoder and query encoder respectively extract information from different views ($\mathcal{H}_{h}$ and $\mathcal{H}_{q}$). As the importance of each view may vary for each node, we cannot simply treat information from the heterogeneous encoder and query encoder equally. Therefore, we introduce a feature fusion module based on the attention mechanism after each layer to take full advantage of complementary information from different views. Specifically, node $v_{i}$ at layer $l$ has two different representations ($\vec{\mathbf{h}}_{i}^{l}$ and $\vec{\mathbf{h}}_{qi}^{l}$), originating from the heterogeneous encoder and the query encoder, respectively. Then, we use an attention mechanism to fuse these complementary representations for node $v_{i}$, as follows:
@@
where $\vec{\mathbf{h}}_{fi}^{l}$ is the fused node representation after $l$ layer, and $\vec{\mathbf{u}}^{\intercal}$ and $\vec{\mathbf{u}}_{q}^{\intercal}$ are the parameterized attention vectors, respectively. $\beta_{i}^{l}$ and $\beta_{qi}^{l}$ can be interpreted as the importance of different views for the node. As shown in Figure 2, for the query encoder, we use the fused node representation as the input for the next layer. For the heterogeneous encoder, we only use the output of the heterogeneous encoder itself in the $l$-th layer as the input of the $(l+1)$-th layer. In this way, we make the heterogeneous encoder independent of the query information, which provides stable prior knowledge about the heterogeneous information networks.
Furthermore, we use the output of the feature fusion module in the final layer as the node representation, which is utilized to compute the probability of each node belonging to a community, formalized as:
(8) $p_{i}=\sigma\Big{(}\text{MLP}(\vec{\mathbf{h}}_{fi}^{L})\Big{)},$
where $p_{i}$ denotes the probability of node $v_{i}$ belonging to the community, $\sigma$ indicates the Sigmoid function, and MLP refers to the multi-layer perceptron.
4.4. Search Algorithm
The search algorithm is designed to identify the target community based on user-specified node types $\mathcal{S}_{\mathcal{A}}$ and the probability of each node belonging to the community calculated by Equation 8. Moreover, in order to ensure the connectivity of communities, we design a search algorithm for multi-type communities in FCS-HGNN based on breadth-first search (BFS).
Input: User-specified node types $\mathcal{S}_{\mathcal{A}}$ in the target community, query node $q$, node probability $p_{i}$, $\mathcal{H}=(V,E,\psi,\phi)$, threshold $\gamma$
@@ 4.4. Search Algorithm
$\mathcal{Q}\leftarrow q$ // Add $q$ into the queue
2 while $\mathcal{Q}$ is not empty do
3       node $v_{i}\leftarrow$ dequeue from $\mathcal{Q}$;
4       if $v_{i}\notin\mathcal{V}$ then
@@
                   $\mathcal{Q}\leftarrow v_{j}$ // add new node for BFS
@@ 7                         $\mathcal{C}_{q}\leftarrow\mathcal{C}_{q}\cup\{v_{j}\}$;
Algorithm 1 Search algorithm for multi-type community
Algorithm 1 outlines our proposed search algorithm for the multi-type community. We first initialize a queue $\mathcal{Q}$ for BFS and a set $\mathcal{V}$ to record visited nodes (Lines 1-3). In the loop, for each neighbor $v_{j}$ of $v_{i}$, we do not restrict the criteria for adding $v_{j}$ to the queue $\mathcal{Q}$, as the neighbors of low probability ($p_{j}<\gamma$) or non-target type nodes ($\psi(v_{j})\notin\mathcal{S}_{\mathcal{A}}$) may also be community members (Line 9). On the contrary, for the node that satisfies the community conditions, we add it to the target community $\mathcal{C}_{q}$ (Lines 10-11). With Algorithm 1, we are able to output the target community based on the probability of each node belonging to the community, while ensuring the connectivity within the community.

### v3

@@ 4. FCS-HGNN
We illustrate the overall framework of FCS-HGNN in Figure 2, which includes three main steps: Heterogeneous and Query Information Preprocessing, Information Integration, and Search Algorithm.
@@ 4.1. Heterogeneous & Query Information Preprocessing
In the community search task over HINs, we primarily have to deal with two aspects of information: the query-dependent information provided by the user and the query-independent information inherent in HINs. Differing from homogeneous graphs, HINs typically comprise multiple types of nodes and edges, presenting significant challenges due to their heterogeneity. Existing works primarily address heterogeneity through predefined meta-paths (Fang et al., 2020; Qiao et al., 2021; Jiang et al., 2022a; Zhou et al., 2023; Liu et al., 2023) or user-specified relationship constraints (Jian et al., 2020), which results in inflexibility and elevates the burden on end-users.
@@ 4.1.1. Type-shared Feature Projection
To address these issues, we propose a type-shared feature projection to deal with the heterogeneity. In HINs, nodes of different types typically reside in distinct feature spaces, which hinders the information aggregation (Wang et al., 2019; Yang et al., 2023). We assume that there exists a unified feature space $\mathcal{U}\in\mathbb{R}^{d_{u}}$. For each node type $t\in\mathcal{T}$, where $\mathcal{T}=\{\psi(v_{i})|v_{i}\in V\}$, we project the original feature of each node $x[v_{i}]$ to the unified feature space $\mathcal{U}$ through a learnable type-shared linear layer:
@@
where $\forall v_{i}\in V$, $t$ represents the type of node $v_{i}$, i.e., $\psi(v_{i})=t$, $W_{t}\in\mathbb{R}^{d_{t}\times d_{u}}$ is a learnable and type-specific linear layer, $d_{u}$ and $d_{t}$ denote the feature dimension of the unified space $\mathcal{U}$ and the original space of node type $t$, respectively. Equation 1 enables nodes of different types to effectively aggregate information in a unified space. It significantly alleviates the burden on systems or end-users as they no longer need to design predefined meta-paths or a series of relationship constraints. It is worth noting that we still preserve the heterogeneity of edges in $\mathcal{H}_{h}$ rather than disregarding it, and propose an edge semantic attention mechanism (elaborated in Section 4.2) to capture more fine-grained heterogeneous information.
@@ 4.1.2. Query-dependent Graph
In addition to the query-independent information inherent in HINs, query information is also very crucial for community search. Therefore, we encode query-dependent information $\mathcal{X}_{q}$ into the HIN $\mathcal{H}$, denoted as $\mathcal{H}_{q}$, where the graph structure maintains consistency with $\mathcal{H}$. Here, $\mathcal{X}_{q}\in\mathbb{R}^{n\times 1}$ is associated with the query node $q$ and is composed of $x_{q}[v_{i}]$, $\forall{v_{i}}\in V$.
@@
The query information $\mathcal{X}_{q}$ indicates whether they are query nodes. The same as $\mathcal{H}_{h}$, we also preserve the type of each edge in $\mathcal{H}_{q}$. This is done to preserve heterogeneous structural information around the query node during the information integration, rather than simply treating $\mathcal{H}_{q}$ as a homogeneous graph.
As shown in Figure 2, we obtain $\mathcal{H}_{h}$ and $\mathcal{H}_{q}$ from $\mathcal{H}$ by employing type-shared feature projection and introducing query-dependent information. In general, $\mathcal{H}_{h}$ and $\mathcal{H}_{q}$ offer two distinct views containing complementary information. The former describes the features of nodes with various types in a unified space, which is beneficial to information extraction and integration in the subsequent steps. The latter delineates the local structural information around the query node, which is crucial for community search.
@@ 4.2.1. Edge Semantic Attention Mechanism
Different from homogeneous graphs, edges in HINs contain complex relationships, which are crucial for understanding the relationships between nodes and mining valuable community information. Therefore, the heterogeneity inherent in edges still requires careful consideration. Existing methods (Fang et al., 2020; Jian et al., 2020; Jiang et al., 2022a; Zhou et al., 2023; Liu et al., 2023) treat different types of relations equally when identifying communities. However, in multi-type communities, the contributions of different relations usually vary across different communities. To address the above issues, we propose the edge semantic attention mechanism as follows:
@@
In real-world HINs, not all networks have edge features. Therefore, we consider the following two scenarios: the HINs include or lack edge features. For the former, we can seamlessly apply equation 3. For the latter, to encode the information of edge types, we introduce an embedding matrix $\mathbf{E}$, which assigns a learnable feature vector to each edge type. We formalize the embedding matrix: $\vec{\mathbf{e}}_{i,j}=\mathbf{E}[t,:]$, where $t\in\{1,2,\ldots,|\mathcal{E}|\}$ is a subscript of =$\mathcal{E}=\{\phi(e_{i,j})|e_{i,j}\in E\}$, indicating the corresponding type of edge $e_{i,j}$, $\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d_{e}}$, $d_{e}$ denotes the feature dimension of edges.
@@ 4.2.2. Heterogeneous Encoder
The heterogeneous encoder is dedicated to capturing query-independent information in $\mathcal{H}_{h}$. Compared to previous studies in HINs (Jian et al., 2020; Fang et al., 2020; Jiang et al., 2022a; Zhou et al., 2023), the heterogeneous encoder dynamically considers the contributions of different relations to communities, rather than treating them equally.
The heterogeneous encoder is equipped with the edge semantic attention mechanism, and iteratively combines the neighbor information as follows:
@@
where $\vec{\mathbf{h}}_{i}^{l}$ denotes the representation of node $v_{i}$ in the $l$-th layer, $\alpha_{ijk}^{l}$ represents the $k$-th edge semantic attention calculated by Equation 3 for selective information aggregation, ||||| | denotes the concatenation, and $\sigma_{g}$ is the ELU (Veličković et al., 2017) activation function. $\mathbf{W}_{r}^{l}\vec{\mathbf{h}}_{i}^{l-1}$ refers to the residual connection designed to prevent over-smoothing and gradient vanishing problems (Li et al., 2018a; Xu et al., 2018). When $l=0$, $\vec{\mathbf{h}}_{i}^{0}=\hat{x}[v_{i}]$.
Through Equation 4, we can uniformly consider structural information and node features with any dimension, avoiding ignoring node features or only considering low-dimensional features as in previous studies (Jian et al., 2020; Fang et al., 2021; Zhou et al., 2023). Compared to ML-based community search methods (Gao et al., 2021; Jiang et al., 2022b; Li et al., 2023) in homogeneous graphs, our proposed heterogeneous encoder integrates edge types and dynamically considers the contributions of different relations instead of treating them equally, which significantly enhances the understanding of complex community patterns in HINs.
4.2.3. Query Encoder
The query encoder is dedicated to obtaining query-dependent information in $\mathcal{H}_{q}$. Compared to previous studies (Jian et al., 2020; Jiang et al., 2022b; Zhou et al., 2023), our proposed query encoder aggregates query information based on different types of edges to promote the model’s understanding of complex relationships in HINs. Specifically, the query encoder consists of multiple GNN layers, each incorporating the edge semantic attention mechanism as follows:
@@ 4.2.3. Query Encoder
where $\vec{\mathbf{h}}_{qi}^{l}$ denotes the query-dependent representation of node $v_{i}$ in the $l$-th layer, and $\vec{\mathbf{h}}_{qi}^{0}=x_{q}[v_{i}]$. Note that for brevity, the learnable parameters, such as $\alpha_{ijk}^{l}$, $\mathbf{W}_{k}^{l}$ and $\mathbf{W_{r}}^{l}$, are no longer distinguished in Equations 4 and 5, but they train these parameters separately. The difference between Equation 4 and Equation 5 lies in their focus on different views, dealing with query-independent and query-dependent information, respectively. Through query encoder, we can effectively capture the local structural information around query nodes to facilitate community search.
@@ 4.2.4. Feature Fusion
The above heterogeneous encoder and query encoder respectively extract information from different views ($\mathcal{H}_{h}$ and $\mathcal{H}_{q}$). As the importance of each view may vary for each node, we cannot simply treat these different information equally. Therefore, we introduce a feature fusion module based on the attention mechanism after each layer to take full advantage of complementary information from different views. Specifically, node $v_{i}$ at layer $l$ has two different representations ($\vec{\mathbf{h}}_{i}^{l}$ and $\vec{\mathbf{h}}_{qi}^{l}$). Then, we use an attention mechanism to fuse these complementary representations for node $v_{i}$, as follows:
@@
where $\vec{\mathbf{h}}_{fi}^{l}$ is the fused node representation after $l$ layer, and $\vec{\mathbf{u}}^{\intercal}$ and $\vec{\mathbf{u}}_{q}^{\intercal}$ are the parameterized attention vectors, respectively. $\beta_{i}^{l}$ and $\beta_{qi}^{l}$ can be interpreted as the importance of different views for the node. As shown in Figure 2, for the query encoder, we use the fused node representation as the input for the next layer. For the heterogeneous encoder, we only use the output of the heterogeneous encoder itself in the $l$-th layer as the input of the $(l+1)$-th layer. In this way, we make the heterogeneous encoder independent of the query information, which provides stable prior knowledge about the HINs.
Furthermore, we use the output of the feature fusion module in the final layer as the node representation, which is utilized to compute the probability of each node belonging to a community:
(7) $p_{i}=\sigma\Big{(}\text{MLP}(\vec{\mathbf{h}}_{fi}^{L})\Big{)}.$
where $\sigma$ indicates the Sigmoid function.
4.3. Depth-based Heuristic Search Algorithm
When dealing with large-scale graphs, traditional community search methods (Jiang et al., 2022a; Liu et al., 2023) usually improve query efficiency by designing appropriate indexes. However, ML-based community search methods (Gao et al., 2021; Jiang et al., 2022b) face significant challenges due to the substantial computational overhead.
To address this issue, we propose the depth-based heuristic search strategy to improve the query efficiency in the large-scale graphs. The core hypothesis is that: the nodes closer to the query node are more likely to belong to the same community. This assumption is grounded in real-life observations: individuals within a community tend to have stronger connections with each other compared to individuals outside the community. In other words, nodes within the same community typically have more edges and shorter paths between them. Therefore, we propose the depth-based heuristic strategy to limit the exploration depth, reducing the search on distant nodes and thereby enhancing search efficiency.
@@ 4.3. Depth-based Heuristic Search Algorithm
$\mathcal{Q}\leftarrow(q,0)$ // Add $q$ and depth 0 into the queue
2
3while $\mathcal{Q}$ is not empty do
4       (node $v_{i}$, depth $d$) $\leftarrow$ dequeue from $\mathcal{Q}$;
5       if $v_{i}\notin\mathcal{V}$ and $d<d_{max}$ then
@@
                   $\mathcal{Q}\leftarrow(v_{j},d+1)$ // add new node for BFS
@@ 8                         $\mathcal{C}_{q}\leftarrow\mathcal{C}_{q}\cup\{v_{j}\}$;
Algorithm 1 Depth-based heuristic search algorithm
Algorithm 1 outlines the depth-based heuristic search algorithm. The search algorithm is designed to identify the target community based on user-specified node types $\mathcal{S}_{\mathcal{A}}$ and the probability of each node belonging to the community calculated by Equation 7. It is worth noting that, for the purpose of subsequent efficiency and performance comparisons, we still employ BFS as the search algorithm for FCS-HGNN. First, we initialize a set $\mathcal{V}$ to record visited nodes, a queue $\mathcal{Q}$ to store node information and record the depth of each node from the query node $q$ (Lines 1-3). Then, we use the depth to determine whether to explore the node (Line 6). In the loop, for each neighbor $v_{j}$ of $v_{i}$, we do not restrict the criteria for adding $v_{j}$ to the queue $\mathcal{Q}$, because the neighbors of low probability ($p_{j}<\gamma$) or non-target type nodes ($\psi(v_{j})\notin\mathcal{S}_{\mathcal{A}}$) may also be community members (Line 9). For the node that satisfies the community conditions, we add it to the target community $\mathcal{C}_{q}$ (Lines 10-11). Through Algorithm 1, we identify target communities based on the node probability while ensuring the connectivity of the communities.